OpenStudy (anonymous):

I have been asked to show that 1-sin 60 deg 1-tan 60 deg ----------- = ----------- cos 60 deg 1+tan 60 deg

OpenStudy (anonymous):

But I am getting LHS = 2 - sqrt(3) but RHS = -2 + sqrt(3)

OpenStudy (anonymous):

Where have I gone wrong??

OpenStudy (lgbasallote):

maybe it's because you substituted the real values instead of proving...

OpenStudy (lgbasallote):

@akitav yours becomes -2 + sqrt 3 too....

OpenStudy (lgbasallote):

@HappySoul maybe they just aren't equal....

OpenStudy (anonymous):

No I did not, I took -2 common in the numerator and cancelled it with the -2 in denominator

OpenStudy (anonymous):

I think you left out a '-' in your original problem statement.

OpenStudy (anonymous):

No, I have double checked the problem....

OpenStudy (anonymous):

screen shot

OpenStudy (anonymous):

you mean (1-sin60)(1-tan60)=cos60(1+tan60) ???

OpenStudy (anonymous):

Not a very good pic but u can see the question.....

OpenStudy (anonymous):

ok wait a minute

OpenStudy (anonymous):

I feel there is some mistake in the question itself.....

OpenStudy (anonymous):

I think it should be \[(\frac{ 1-\sin 60^{0} }{ \cos60^{0}}) equals -(\frac{ 1-\tan60^{0} }{ 1+\tan60^{0} }) \]

OpenStudy (anonymous):

damn.... i thought what is the problem..

OpenStudy (anonymous):

is my thinking as above correct??

OpenStudy (anonymous):

try by doing cross multiplication

OpenStudy (anonymous):

it is not calculate in number.. its purpose is using trigonometrical function eqation.

OpenStudy (anonymous):

try convert tan to sin and cos.

OpenStudy (anonymous):

Since a specific angle is given, we have to use the value of the angle here (as per examples in the book)....... Pls try and solve by the method suggested by you, if you can get the answer, pls post it...thanks.....☺

OpenStudy (anonymous):

ok i am doing it.

OpenStudy (anonymous):

may be question is wronng pleasse check it twice.

OpenStudy (anonymous):

yeah.. i am really stuck in proving

OpenStudy (anonymous):

\[\frac{ 1-\sin60 }{ \cos60 }=\frac{ \cos60 - \sin60 }{ \cos60 + \sin60 } \]

OpenStudy (anonymous):

what can i do later...